OpenStudy (anonymous):

sole x^2 - 12x - 13 = 0

OpenStudy (dumbcow):

find factors of -13 that add up to -12 -13 and 1 --> (x-13)(x+1) = 0 x-13 = 0 x = 13 x+1 = 0 x = -1